# Geometry-Aware Edge Pooling for Graph Neural Networks

### v1

@@ 3.2 Theoretical Analysis
We now present theoretical properties of our pooling methods. First, we highlight fundamental invariances and properties of magnitude and spread used to design our algorithms. Further, we provide a bound on the difference in magnitude during pooling by the difference in spread, demonstrating the close relationship between the two. For a complete list or theorems and proofs see Section B.3.
@@
Consider a positive definite metric graph $G=G_{1}\sqcup G_{2}$ consisting of the disjoint union of two graphs $G_{1}$ and $G_{2}$. Then $\text{Mag}(G)=\text{Mag}(G_{1})+\text{Mag}(G_{2})$.
@@
Our pooling layers are invariant under isometries of the input graph provided that the choice of edges to contract at each iteration is deterministic in the case that the edge scores coincide. This is a consequence of the following result.
@@
If the metric graphs $G_{1}$ and $G_{2}$ are isometric then $\text{Mag}(G_{1})=\text{Mag}(G_{2})$ and $\text{Sp}(G_{1})=\text{Sp}(G_{2})$.
@@
Edge contraction is the main operation of our pooling methods and can be considered as a map $f$ between graphs. The following result provides a sufficient condition to ensure that the morphism $f$ is compatible with the metric structure.
@@
Consider an edge-contraction map $f:(G_{1},d_{1})\rightarrow(G_{2},d_{2})$ between positive definite metric graphs. If the map is Lipschitz (i.e. $d_{2}(f(v_{1}),f(v_{2}))\leq d_{1}(v_{1},v_{2})\ \forall v_{1},v_{2}\in X_{1}$) then $\text{Mag}(G_{2})\leq\text{Mag}(G_{1})$.
Starting from an initial metric graph $G=(X,E)$, a Lipschitz edge-contraction map $f$ yields a sequence of graphs $\{G_{i}\}_{i=1}^{k}$ where $k$ is the number of edges that have been contracted. This sequence can be constructed as described by our algorithms in Section 3.1. Note that it is not guaranteed that MagEdgePool and SpreadEdgePool yield the same sequence of hierarchically pooled graphs. For this reason, we will refer to the graphs resulting from the $k^{th}$ edge-contraction with magnitude and spread as $G^{(k)}$ and $\widetilde{G}^{(k)}$, respectively. Let $\Delta^{(k)}\text{Mag}(G)=|\text{Mag}(G^{(k-1)})-\text{Mag}(G^{(k)})|$ and let $\Delta^{(k)}\text{Sp}(G)=|\text{Sp}(G^{(k-1)})-\text{Sp}(G^{(k)})|$.
@@
The time complexity of our pooling methods is independent of the GNN and is dominated by the cost of computing the edge scores in Equation 5. Given a graph $G=(X,E)$, magnitude has time complexity $O(|X|^{3})$. In comparison, spread has time complexity $O(|X|^{2})$ and can be more efficiently approximated [18]. Spread thus offers a considerably faster alternative. Now, let $O(C_{d})$ be the time complexity of computing the metric on $G$. The time complexity of SpreadEdgePool is dominated by $O(|E|(C_{d}+|X|^{2}+\text{log}|E|))$ and MagEdgePool by $O(|E|(C_{d}+|X|^{3}+\text{log}|E|))$. We note that on large graphs, it is possible to speed up the distance computations further to ensure scalability. See Section B.2 for a full description of computational costs and Section D.2 for an empirical evaluation, which shows that our algorithm performs on-par with existing pooling methods across the datasets evaluated throughout this work.
@@ 4 Experimental Results
Across our experiments, we address the following three key tasks: (i) graph classification performance, (ii) graph structure preservation during pooling, (iii) and performance across varying pooling ratios.
@@ 4.1 Graph Classification
Table 1: Mean and standard deviation of the graph classification accuracy. The best performing model is marked in bold. All models that did not perform significantly different from the best model are coloured green. The rightmost column shows the mean rank of each pooling method across datasets.
@@
A key aim of implementing graph pooling layers is to preserve task-relevant information while reducing computational costs. In particular, useful pooling layers guarantee good performance across a wide range of different datasets, thus capturing essential information for the task at hand. We thus investigate how well our aim to preserve structural diversity during edge pooling translates to practical performance at graph classification tasks.
Experimental Setup.
We evaluate 8 different graph datasets [35], as detailed in Section C.2. Whenever node features are not available, we use node degree as an input feature. To ensure a fair comparison with alternative pooling methods, we follow the experimental setup by Grattarola et al. [25] and guidance by Errica et al. [19] for model selection. Specifically, we plug in each pooling layer into the model architecture specified by Grattarola et al. [25], which is of the following form:
The model includes pre-processing and post-processing MLPs with 2 layers, 256 hidden units, ReLU activation, and batch normalization. $\text{GNN}(\mathbf{X},\mathbf{A})$ refers to a graph neural network layer, more specifically a general convolutional layer [46] with parameters chosen according to the best results achieved by You et al. [46]. As an intermediate layer, $\text{POOL}(\mathbf{X},\mathbf{A})$ corresponds to a specific pooling layer. All pooling layers are configured to pool each graph to around 50% of nodes. We also compare with ‘No Pooling,’ the same model architecture without any pooling layers. We use 10-fold stratified cross-validation and further partition the training data into 90% training and 10% validation data while keeping the labels balanced between splits. Finally, we report the best test accuracy of each model trained using Adam with a cross-entropy loss (batch size 32, learning rate $0.0005$, and early stopping based on the validation loss with a patience of 50 epochs). Further details are described in Section C.5.2.
@@
Table 1 reports the mean and standard deviation of the test accuracy achieved by different pooling methods. We furthermore highlight which methods do not perform statistically significantly different from the best model (using pairwise Wilcoxon signed-rank tests applied to the accuracy scores and employing Holm–Bonferroni correction at a significance threshold of $p=0.05$), thus permitting us to identify pooling methods that achieve top performance. Notably, both MagEdgePool and SpreadEdgePool achieve the best mean ranks across datasets in terms of their accuracy. Further, they are always among the top-performing methods across all evaluated datasets. Altogether, both their ranking and their individual accuracy scores thus demonstrate superior and consistently high performance across graph classification tasks. The performance benefits of our proposed pooling methods are most pronounced on DHFR [40], where our methods surpass even the GNN without pooling layer by around 17 percentage points. This provides evidence that the regularising effects of our pooling approach can help reduce overfitting, especially for small datasets and geometrically-rich graphs. For other biological datasets (Mutagenicity, NCI1 and NC109), our methods show competitive performance with trainable layers, indicating that the introduction of additional trainable components into the pooling layer is not necessary to guarantee high task performance. On ENZYMES, DHFR, IMDB-BINARY, and IMDB-MULTI, non-trainable methods generally outperform trainable pooling layers, with MagEdgePool and SpreadEdgePool consistently reaching high accuracy.
Comparing pooling methods to using no pooling, we observe that MagEdgePool and SpreadEdgePool reach similar or even higher performance across datasets. For six datasets, our diversity-guided pooling methods improve mean accuracy, indicating that pooling retains task-relevant information while aiding the generalisation capabilities of the GNN. MagEdgePool and SpreadEdgePool act as interpretable and expressive pooling transformations, which reduce the computational costs making GNNs learn from graphs’ coarsened geometry. As reported in Table 1, MagEdgePool and SpreadEdgePool achieve very similar accuracies across datasets superior to alternative pooling layers. In practice, especially for large graphs, we recommend using SpreadEdgePool due to its high predictive performance and superior computational efficiency.
@@ 4.2 Magnitude and Graph Structure Preservation
Figure 3: Structure preservation for all graphs in the NCI1 dataset across varying pooling ratios. Left: The spectral distance between the normalised Laplacians of the the original and the pooled graphs. Right: The relative difference in magnitude, which summarises the proportional difference in structural diversity after pooling. Violinplots show the variability across graphs at pooling ratio 0.5.
@@
Motivated by the visual comparison of graphs pooled using different pooling layers from Figure 1, we next set out to investigate the link between structure preservation and task performance. Specifically, we choose NCI1, a dataset of 4,110 graphs corresponding to chemical compounds, because it has been shown to possess both informative features and task-relevant graph structures [13]. We follow the same classification procedure as before and extract the pooled graph representations after training the GNNs described in Table 1. Three pooling layers, MinCut, DiffPool and NMF, were removed from further comparison because they showed notably worse qualitative results for the motivating examples in Figure 1 and classification performance in Table 1. NDP and Graclus are evaluated across fewer pooling ratios than more adaptive methods, because they pool graphs to around half their size at every step. To assess graph structure preservation we use the spectral distance defined as $\sqrt{\sum_{k=1}^{K}(\lambda_{k}-\lambda_{k}^{\prime})^{2}}$, the $l_{2}$-norm between the eigenspectra of the normalised Laplacians of the original and the pooled graphs [42]. We also report the magnitude difference between graphs to evaluate the preservation of structural diversity.
@@
MagEdgePool and SpreadEdgePool show low spectral distances across pooling ratios as visualised in Figure 3. This indicates that contracting edges guided by structural diversity preserves key spectral properties. Our methods also demonstrate low magnitude differences confirming that they perform as intended. In fact, the structure preservation scores for MagEdgePool and SpreadEdgePool coincide almost perfectly, giving empirical evidence that spread offers an alterative to magnitude. Figure 3 indicates that preserving magnitude corresponds to lower spectral distances and better retention of spectral properties. This link supports our motivation of guiding pooling by magnitude.
@@
Alternative pooling layers fail to effectively preserve graphs’ structural properties during both qualitative and quantitative comparisons to varying extents. Node decimal pooling (NDP) [9] was specifically designed to preserve spectral properties during pooling. However, it still reaches both higher spectral distances and higher magnitude differences than MagEdgePool on average across pooling ratios, as visualised in Figure 3. Finally, the sparse pooling layers Graclus, TopKPool, and SAGPool, all show higher spectral distances than our approach. This difference is even more pronounced in terms of magnitude differences, where all these three methods demonstrate high distortion of the underlying metric space diversity. These findings agree with the qualitative comparisons between graphs pooled using different pooling layers as visualised in Figure 1. We thus conclude that MagEdgePool and SpreadEdgePool successfully encode graphs’ coarsened geometry during pooling, surpassing alternative pooling methods.
@@ 4.3 Pooling Ratio and Task Performance
From Table 1 we observe that it is possible to reach very high performance on benchmark datasets even while pooling each graph to half its size. Based on this, we further investigate how the pooling ratio influences pooling layers and their classification performance. We consider two datasets, NCI1 and ENZYMES, which contains 600 graphs representing protein tertiary structures from 6 classes of enzymes and is selected as an example of a multi-class prediction task. We keep the same experimental setup described in Table 1, but use GIN layers instead of general convolutional layers to further assess whether the trends in performance differ across model architectures. Figure 4 then reports the classification accuracy achieved by each pooling layer for varying pooling ratios.
@@
Notably, we observe that MagEdgePool and SpreadEdgePool consistently reach very high test accuracies even at low pooling rations. Meanwhile, the performance of other non-trainable methods drops notably more when graphs are pooled to up to 6.25% of their original size showing that they fail to preserve task-relevant information for both ENZYMES and NCI1. We note that the trainable pooling layers, TopK and SAGPool, reach higher or comparable performance on NCI1 for pooling ratios above 50%, but decrease in accuracy for more extreme pooling ratios. They consistently perform worse on ENZYMES, indicating that they distort important graph features or key graph structures during pooling. MagEdgePool and SpreadEdgePool in comparison reach superior performance and lower decreases in accuracy across varying pooling ratios demonstrating their potential to offer reliable, interpretable and stable pooling operations. Overall, the reported accuracies in Figure 4 agree with results in Table 1 indicating that our observations hold for varying choices of GNN layers. We thus find across experiments that MagEdgePool and SpreadEdgePool constitute useful general-purpose pooling approaches, demonstrating their capability to faithfully encoding graphs’ geometry, which ensures stable performance across pooling ratios, datasets and GNN architectures.

### v2

@@ 3.2 Theoretical Analysis
We now present theoretical properties of our pooling methods. First, we highlight fundamental invariances and properties of magnitude and spread used to design our algorithms. Further, we provide a bound on the difference in magnitude during pooling by the difference in spread, demonstrating the close relationship between the two. For a complete list or theorems and proofs, please refer to Section˜B.3.
@@
Consider a graph $G=G_{1}\sqcup G_{2}$ consisting of the disjoint union of two graphs $G_{1}$ and $G_{2}$. Then $\text{Mag}(G)=\text{Mag}(G_{1})+\text{Mag}(G_{2})$.
@@
Our pooling layers are invariant under isometries of the input graph provided that the edge choice at each iteration is deterministic whenever the edge scores coincide. This is a consequence of the following result.
@@
For isomorphic graphs $G_{1},G_{2}$, we have $\text{Mag}(G_{1})=\text{Mag}(G_{2})$ and $\text{Sp}(G_{1})=\text{Sp}(G_{2})$.
@@
Edge contraction is the main operation of our pooling methods and can be considered as a map $f$ between graphs. The following result provides a sufficient condition to ensure that $f$ remains compatible with the metric structure.
@@
Consider an edge-contraction map $f\colon(G_{1},d_{1})\rightarrow(G_{2},d_{2})$ between positive definite metric graphs. If the map is $1$-Lipschitz, i.e. $d_{2}(f(v_{1}),f(v_{2}))\leq d_{1}(v_{1},v_{2})\ \forall v_{1},v_{2}\in X_{1}$, then $\text{Mag}(G_{2})\leq\text{Mag}(G_{1})$.
Starting from an initial metric graph $G=(X,E)$, a 1-Lipschitz edge-contraction map $f$ yields a sequence of graphs $\{G_{i}\}_{i=1}^{k}$, where $k$ is the number of edges that have been contracted. This sequence can be constructed as described by our algorithms in Section˜3.1. We refer to the graphs resulting from the $k^{th}$ edge-contraction using MagEdgePool as $G^{(k)}$. Let $\Delta^{(k)}\text{Mag}(G)=|\text{Mag}(G^{(k-1)})-\text{Mag}(G^{(k)})|$ and let $\Delta^{(k)}\text{Sp}(G)=|\text{Sp}(G^{(k-1)})-\text{Sp}(G^{(k)})|$.
@@
Expressivity.
Studying the expressive power of GNNs and their ability for distinguishing non-isomorphic graphs offers theoretical insights for understanding the theoretical capabilities of pooling operators. Bianchi and Lachi [8] state sufficient condition for a pooling layer to preserve the expressive power of the preceding message-passing (MP) layers. As demonstrated in Section˜B.4, MagEdgePool and SpreadEdgePool satisfy these conditions ensuring expressivity.
@@
The time complexity of our pooling methods is independent of the GNN and is dominated by the cost of computing the edge scores in Equation˜5. Given a graph $G=(X,E)$, magnitude has time complexity $O(|X|^{3})$. In comparison, spread has time complexity $O(|X|^{2})$ and can be more efficiently approximated [23]. Spread thus offers a considerably faster alternative. Now, let $O(C_{d})$ be the time complexity of computing the metric on $G$. The time complexity of SpreadEdgePool is dominated by $O(|E|(C_{d}+|X|^{2}+\text{log}|E|))$ and MagEdgePool by $O(|E|(C_{d}+|X|^{3}+\text{log}|E|))$. We note that on large graphs, it is possible to speed up the distance computations further to ensure scalability. See Section˜B.2 for a full description of computational costs and Section˜D.2 for an empirical evaluation, which shows that our algorithm performs on-par with existing pooling methods across the datasets evaluated throughout this work.
@@ 4 Experimental Results
Across our experiments, we address four key tasks, namely (i) graph classification performance, (ii) graph structure preservation during pooling, (iii) performance across varying pooling ratios, and (iv) performance at graph property regression.
@@ 4.1 Graph Classification
The primary aim of using graph pooling layers is to preserve task-relevant information while reducing computational costs. In particular, useful pooling layers guarantee good performance across a wide range of different datasets, thus capturing essential information for the task at hand. We thus investigate how well our aim to preserve structural diversity during edge pooling translates to practical performance at graph classification tasks. Note that our goal is not to reach state-of-the-art accuracies on all tasks, but to benchmark the performance gain or loss of different pooling operators.
Experimental Setup.
We evaluate 8 different graph datasets, as detailed in Section˜C.2. Whenever node features are not available, we use node degree as an input feature. To ensure a fair comparison with alternative pooling methods, we follow the experimental setup by Grattarola et al. [32] and guidance by Errica et al. [24] for fair model comparison. Specifically, we plug in each pooling layer into the model architecture specified by Grattarola et al. [32], which is of the following form:
The model includes pre-processing and post-processing MLPs with 2 layers, 256 hidden units, ReLU activation, and batch normalization. $\text{GNN}(\mathbf{X},\mathbf{A})$ refers to a graph neural network layer, more specifically a general convolutional layer [57] with parameters chosen according to the best results achieved by You et al. [57]. As an intermediate layer, $\text{POOL}(\mathbf{X},\mathbf{A})$ corresponds to a specific pooling layer. All pooling layers are configured to pool each graph to around 50% of nodes. We also compare with ‘No Pooling,’ the same model architecture without any pooling layers. We use 10-fold stratified cross-validation and further partition the training data into 90% training and 10% validation data while keeping the labels balanced between splits. Finally, we report the best test accuracy of each model trained using Adam with a cross-entropy loss (batch size 32, learning rate $0.0005$, and early stopping based on the validation loss with a patience of 50 epochs). Further details are described in Section˜C.5.3.
Table 1: Mean and standard deviation of the graph classification accuracy. The best-performing model is marked in bold. All models that did not perform significantly different from the best model are coloured green. The rightmost column shows the mean rank of each pooling method across datasets.
@@
Table˜1 reports the mean and standard deviation of the test accuracy achieved by different pooling methods. We furthermore highlight which methods do not perform statistically significantly different from the best model (using pairwise Wilcoxon signed-rank tests applied to the accuracy scores and employing Holm–Bonferroni correction at a significance threshold of $p=0.05$), thus permitting us to identify pooling methods that achieve top performance. Notably, both MagEdgePool and SpreadEdgePool achieve the best mean ranks across datasets in terms of their accuracy. Further, they are always among the top-performing methods across all evaluated datasets. Altogether, both their ranking and their individual accuracy scores thus demonstrate superior and consistently high performance across graph classification tasks. The performance benefits of our proposed pooling methods are most pronounced on DHFR [49], where our methods surpass even the GNN without pooling layer by around 17 percentage points. This provides evidence that the regularising effects of our pooling approach can help reduce overfitting, especially for small datasets and geometrically-rich graphs. For other biological datasets (Mutagenicity, NCI1 and NC109), our methods show competitive performance with trainable layers, indicating that the introduction of additional trainable components into the pooling layer is not necessary to guarantee high task performance. On ENZYMES, DHFR, IMDB-BINARY, and IMDB-MULTI, non-trainable methods generally outperform trainable pooling layers, with MagEdgePool and SpreadEdgePool consistently reaching high accuracy.
Comparing pooling methods to using no pooling, we observe that MagEdgePool and SpreadEdgePool reach similar or even higher performance across datasets. For six datasets, our diversity-guided pooling methods improve mean accuracy, indicating that pooling retains task-relevant information while aiding the generalisation capabilities of the GNN. MagEdgePool and SpreadEdgePool act as interpretable and expressive pooling transformations (see Section˜D.3) , which reduce the computational costs making GNNs learn from graphs’ coarsened geometry. As reported in Table˜1, MagEdgePool and SpreadEdgePool achieve very similar accuracies across datasets reaching top accuracies compared to alternative pooling layers. In practice, especially for large graphs, we recommend using SpreadEdgePool due to its high predictive performance and superior computational efficiency.
@@ 4.2 Magnitude and Graph Structure Preservation
Motivated by the visual comparison of graphs pooled using different pooling layers from Figure˜1, we next set out to investigate the link between structure preservation and task performance. Specifically, we choose NCI1, a dataset of 4,110 graphs corresponding to chemical compounds, because it has been shown to possess both informative features and task-relevant graph structures [15]. We follow the same classification procedure as before and extract the pooled graph representations after training the GNNs described in Section˜4.1. Three pooling layers, MinCut, DiffPool and NMF, were removed from further comparison because they showed notably worse qualitative results for the motivating examples in Figure˜1 and classification performance in Table˜1. NDP and Graclus are evaluated across fewer pooling ratios than more adaptive methods, because they pool graphs to around half their size at every step. To assess graph structure preservation we use the spectral distance defined as $\sqrt{\sum_{k=1}^{K}(\lambda_{k}-\lambda_{k}^{\prime})^{2}}$, i.e. the $l_{2}$-norm between the eigenspectra of the normalised Laplacians of the original and the pooled graphs [53]. We also report the magnitude difference between graphs to evaluate the preservation of structural diversity.
@@
MagEdgePool and SpreadEdgePool exhibit small spectral distances across pooling ratios as visualised in Figure˜3. This indicates that contracting edges guided by structural diversity preserves key spectral properties. Our methods also demonstrate low magnitude differences confirming that they perform as intended. In fact, the structure preservation scores for MagEdgePool and SpreadEdgePool coincide almost perfectly, giving empirical evidence that spread offers an alterative to magnitude. Figure˜3 indicates that preserving magnitude corresponds to lower spectral distances and better retention of spectral properties. This link supports our motivation of guiding pooling by magnitude.
@@
Alternative pooling layers fail to effectively preserve graphs’ structural properties during both qualitative and quantitative comparisons to varying extents. Node decimal pooling (NDP) [10] was specifically designed to preserve spectral properties during pooling. However, it still reaches both higher spectral distances and higher magnitude differences than MagEdgePool on average across pooling ratios, as visualised in Figure˜3. Finally, the sparse pooling layers Graclus, TopKPool, and SAGPool, all show higher spectral distances than our approach. This difference is even more pronounced in terms of magnitude differences, where all these three methods demonstrate high distortion of the underlying metric space diversity. These findings are repeated for further datasets (see Figure˜S.7 in the appendix) and agree with the qualitative comparisons between graphs pooled using different pooling layers as visualised in Figure˜1. We thus conclude that MagEdgePool and SpreadEdgePool successfully encode graphs’ coarsened geometry during pooling, surpassing alternative pooling methods.
Figure 3: Structure preservation for all graphs in the NCI1 dataset across varying pooling ratios. Left: The spectral distance between the normalised Laplacians of the original and the pooled graphs. Right: The relative difference in magnitude, which summarises the proportional difference in structural diversity after pooling. Violin plots show the variability across graphs at pooling ratio 0.5.
@@ 4.3 Pooling Ratio and Task Performance
From Table˜1, we observe that it is possible to reach very high performance on benchmark datasets even while pooling each graph to half its size. Based on this, we further investigate how the pooling ratio influences pooling layers and their classification performance. We consider two datasets, NCI1 and ENZYMES, which contains 600 graphs representing protein tertiary structures from 6 classes of enzymes and is selected as an example of a multi-class prediction task. We keep the same experimental setup described in Section˜4.1, but use GIN layers instead of general convolutional layers to further assess whether the trends in performance differ across model architectures. Figure˜4 reports the accuracy achieved by each pooling layer for varying pooling ratios. Notably, we observe that MagEdgePool and SpreadEdgePool consistently reach very high test accuracies even at low pooling rations. Meanwhile, the performance of other non-trainable methods drops notably more when graphs are pooled to up to 6.25% of their original size showing that they fail to preserve task-relevant information for both ENZYMES and NCI1. We note that the trainable pooling layers, TopK and SAGPool, reach higher or comparable performance on NCI1 for pooling ratios above 50%, but decrease in accuracy for more extreme pooling ratios. They consistently perform worse on ENZYMES, indicating that they distort important graph features or key graph structures during pooling. MagEdgePool and SpreadEdgePool in comparison reach top performance and lower decreases in accuracy across varying pooling ratios demonstrating their potential to offer reliable, interpretable and stable pooling operations. Overall, the reported accuracies in Figure˜4 agree with results in Table˜1 indicating that our observations hold for varying choices of GNN layers. We thus find across experiments that MagEdgePool and SpreadEdgePool constitute useful general-purpose pooling approaches, demonstrating their capability to faithfully encoding graphs’ geometry, which ensures stable performance across pooling ratios, datasets and GNN architectures.
Table 2: RMSE on the test data for different pooling methods.
Method MolEsol MolFreesolv MolLipo
No Pooling 1.44 ± 0.10 2.98 ± 0.78 0.98 ± 0.30
MagEdge 1.47 ± 0.11 2.81 ± 0.30 0.91 ± 0.23
SpreadEdge 1.58 ± 0.15 2.83 ± 0.24 0.91 ± 0.20
NDP 1.54 ± 0.18 3.17 ± 0.29 0.82 ± 0.09
Graclus 1.47 ± 0.13 2.99 ± 0.27 0.85 ± 0.17
NMF 2.37 ± 0.77 15.06 ± 8.83 0.85 ± 0.04
TopK 1.68 ± 0.10 3.25 ± 0.54 1.07 ± 0.25
SAGPool 1.66 ± 0.18 2.77 ± 0.18 1.03 ± 0.20
DiffPool 1.71 ± 0.14 5.78 ± 4.55 2.35 ± 2.57
MinCut 1.88 ± 0.51 4.28 ± 0.32 1.22 ± 0.19
4.4 Graph Regression
We further apply our pooling methods to graph regression tasks. We modify the GNN architecture as described in Section˜C.5.4 and plug in varying pooling layers. Table˜2 reports the RMSE on the test dataset across pre-defined data splits and ten random seeds for three datasets from OGB [33]. Overall, these results confirm that SpreadEdgePool and MagEdgePool constitute useful pooling operations and reach comparatively low RMSEs across these three regression tasks. Our pooling methods successfully retain or improve task performance, indicating their ability to preserve task-relevant information.
4.5 Computational Efficiency
Finally, we evaluate the computational efficiency and scalability of pooling operators (see Section˜D.2 for details). Our approaches scale reasonably well to the datasets evaluated in this study, both in terms of runtime and memory requirements, thus remaining applicable for graphs with up to few hundreds of nodes. Experiments confirm that SpreadEdgePool is notably faster than MagEdgePool and that distance approximations can be used to speed up the edge score computations used for pooling. Further, we find that pre-computing allows for notably lower GNN training costs compared to trainable methods; when comparing our methods to EdgePool, our methods exhibit superior computational efficiency during training while ensuring similarly high classification performance.